OpenStudy (anonymous):

I WILL MEDAL CX A stack of quarters is packed in a plastic tube 4 inches tall. The diameter of each quarter is 0.94 inch. What is the approximate surface area of the cylinder formed by the stack of coins? (Use 3.14 for .) 2.77 in2 18.4 in2 3.76 in2 13.19 in2

OpenStudy (anonymous):

Sa=2πrh+2πr2 h= 4 r=.94/2 = .47 sa=2π(.47)(4)+2π(.47)2 sa=2π(1.88)+2π(.2209) sa=3.76π+.4418π sa=4.2018π sa=4.2018×3.14 sa=13.193652≈13.19

OpenStudy (anonymous):

whoaa my head is gonna explode XD thanks cx

OpenStudy (anonymous):

no prob